### v1

# Region-of-Interest-Guided Deep Joint Source-Channel Coding for Image Transmission

Abstract
Deep joint source-channel coding (deepJSCC) and semantic communication have shown promising improvements in communication performance over wireless networks. However, current approaches primarily focus on enhancing average performance metrics, such as overall image reconstruction quality or task accuracy, which may not fully align with users’ actual experience—often driven by the quality of specific regions of interest (ROI). Motivated by this, we propose ROI-guided joint source-channel coding (ROI-JSCC), a novel deepJSCC framework that prioritizes high-quality transmission of ROI. The ROI-JSCC consists of four key components: (1) ROI embedding and feature map extraction, (2) ROI-guided split processing, (3) ROI-based loss function design, and (4) ROI-adaptive bandwidth allocation. Together, these components enable ROI-JSCC to selectively improve the reconstruction quality of varying ROI while preserving overall image quality without increasing computational burden. Experimental results under diverse communication environments demonstrate that ROI-JSCC significantly improves ROI reconstruction quality while maintaining competitive average image quality compared to recent state-of-the-art methods. All codes are available at https://github.com/hansung-choi/ROI-JSCC.
Index Terms:
joint source-channel coding, region-of-interest, image transmission, quality of experience
I Introduction
With the arrival of the 6G era, a wide range of applications is anticipated, including autonomous driving, smart factories, digital twins, and extended reality technologies such as virtual and augmented reality (VR/AR) [1]. These applications place immense demands on wireless networks, particularly in terms of reliability and latency. However, existing digital communication systems—refined over decades to ensure bit-level accuracy—struggle to meet these requirements. Further improvements to conventional systems have become increasingly difficult, posing significant challenges in supporting the demanding transmission needs of next-generation applications.
To address these challenges, deep learning-based joint source-channel coding (deepJSCC) methods have been proposed for various data modalities, including images [2, 3], point cloud [4], text [5], and speech [6]. These approaches leverage the capability of deep learning models to jointly learn the characteristics of source data and communication channels, enabling the mapping of source data directly into robust and informative channel input symbols. In contrast to traditional communication systems that often fail under severe channel conditions (e.g., high noise levels), deepJSCC methods are able to reliably transmit the meaning of the data. Furthermore, its integration into downstream tasks has led to the development of task-oriented semantic communication (SemCom) systems. By transmitting only task-relevant information, SemCom significantly reduces communication overhead while maintaining high downstream task performance in areas such as image segmentation [7], super-resolution [8], and image classification [9].
Although deepJSCC and SemCom methods have shown great success in preserving the average meaning of data and improving overall task performance, few approaches explicitly consider the ultimate goal of communication in emerging applications. For instance, in autonomous driving, maximizing average semantic segmentation accuracy is not the ultimate objective. Instead, prioritizing performance in important regions, such as pedestrians nearby, even at the cost of reduced segmentation accuracy for less relevant regions like the sky or distant static objects, may contribute more to safe and efficient driving. Similarly, in extended reality, reducing the average reconstruction error between the original and reconstructed images may have a limited impact on the user’s experience. The human visual system is highly sensitive to details in the fovea region, the area near the gaze point, while its sensitivity drops sharply in peripheral regions, the area distant from the gaze point [10]. Therefore, minimizing distortion in the user’s fovea region, even at the expense of decreased quality in peripheral can enhance perceived visual quality. Nevertheless, current deepJSCC and SemCom frameworks lack appropriate coding mechanisms that account for these functional priorities for human experiences.
To address the above issue, we propose a region-of-interest-guided deep joint source-channel coding (ROI-JSCC) for user experience-oriented image transmission. Through the following four mechanisms, the ROI-JSCC effectively preserves high quality in regions of interest (ROI), which dominates user experience: (1) ROI embedding and feature map extraction: This mechanism integrates ROI information into intermediate deepJSCC features, allowing the model to focus on preserving ROI-specific content. (2) ROI-guided split processing: Important ROI features are refined using intensive self-attention [3], while less important features are processed with lightweight spatial attention [11]. This enhances the processing capability for ROI features without incurring significant computational overhead. (3) ROI-based loss function design: Leveraging the mechanism in (1), this loss function guides ROI-JSCC to learn which features are related to ROI and how to encode them to preserve their quality in the transmitted image. (4) ROI-adaptive bandwidth allocation: This mechanism significantly increases bandwidth allocation for ROI features while slightly reducing it for others, thereby enhancing ROI quality with minimal degradation elsewhere.
These mechanisms enable ROI-JSCC to improve ROI-specific image quality across varying ROI configurations without requiring fine-tuning. Moreover, it efficiently utilizes computational resources. Experimental results demonstrate that ROI-JSCC significantly enhances ROI quality while maintaining competitive computational efficiency compared to existing deepJSCC approaches [12, 3, 13, 14].
Figure 1: ROI-JSCC communication system.
II ROI-JSCC Communication System
We assume that an image is partitioned into $n_{h}\times n_{w}$ patches. The patches can then be classified into three categories based on their impact on user experience: region of interest (ROI), region of periphery (ROP), and region of non-interest (RONI), as depicted in Figure 4, the ROI map. In this map, ROPs denote patches adjacent to the ROI, while RONIs represent all remaining regions. The importance of a patch decreases as its distance from the ROI increases. We further assume that there is only one ROI and the position of ROI $\gamma:=(h_{\gamma},w_{\gamma})$ in the image is shared between transmitter and receiver, where $h_{\gamma}$ and $w_{\gamma}$ denote the vertical and horizontal coordinates of the ROI in the ROI map, respectively. Our objective is to transmit images in a manner that preserves high quality in ROI, which dominates user experience, while minimizing degradation in overall image quality. To this end, we propose a ROI-JSCC communication system illustrated in Figure 1.
In this system, the encoder $f_{\theta}$ extracts a semantic feature matrix $\mathbf{Z}:=[\mathbf{z}_{1},\ldots,\mathbf{z}_{B}]^{T}\in\mathbb{C}^{B\times C_{m}}$ from a source image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ using the ROI position $\gamma$, where $H$, $W$, and $3$ denote the height, width, and RGB channels of $\mathbf{x}$, respectively. Among the semantic features, some contain significantly more information about the ROI; we refer to these as ROI feature vectors, denoted by $\mathbf{z}_{{\textrm{ROI}}}$. To preserve the information in $\mathbf{z}_{{\textrm{ROI}}}$, the transmitter allocates greater bandwidth to $\mathbf{z}_{{\textrm{ROI}}}$ and less bandwidth to the remaining features by transforming each feature vector dimension, under a total bandwidth constraint. The feature vectors are then aggregated into $\mathbf{\tilde{z}}\in\mathbb{C}^{k}$, where $k$ denotes the total channel bandwidth usage. The channel usage per pixel (CPP) is defined as the ratio of the channel bandwidth to the total number of RGB pixels in $\mathbf{x}$, i.e., ${\textrm{CPP}}:=\frac{k}{3HW}$. Next, the transmitter rescales $\mathbf{\tilde{z}}$ into $\mathbf{\bar{z}}$ to fit the power constraint $P$. Throughout this paper, we assume $P=1$, i.e., $\frac{1}{k}\left\lVert\mathbf{\bar{z}}\right\rVert_{2}^{2}\leq 1$.
After power normalization, $\mathbf{\bar{z}}$ is transmitted over a wireless noisy channel. The receiver then obtains the corrupted feature vector $\mathbf{\hat{z}}_{r}\in\mathbb{C}^{k}$. In this paper, we consider an additive white Gaussian noise (AWGN) channel, modeled as
$\mathbf{\hat{z}}_{r}=\mathbf{\bar{z}}+\mathbf{\epsilon},~{}~{}\mathbf{\epsilon}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{k})$
where $\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{k})$ denotes a circularly symmetric complex Gaussian distribution with zero mean and covariance matrix $\sigma^{2}\mathbf{I}_{k}$. The quality of the AWGN channel is characterized by the signal-to-noise ratio (SNR), defined as
${\textrm{SNR}}:=10\log_{10}\left(1/\sigma^{2}\right)~{}\text{dB}.$
At the receiver, the received vector $\mathbf{\hat{z}}_{r}\in\mathbb{C}^{k}$ is recovered into $\mathbf{\hat{Z}}^{p}:=[\mathbf{\hat{z}}^{p}_{1},\ldots,\mathbf{\hat{z}}^{p}_{B}]^{T}\in\mathbb{C}^{B\times C_{m}}$ using zero-padding based on the ROI position $\gamma$, as illustrated in Figure 1. The decoder $g_{\phi}$ then reconstructs an image $\mathbf{\hat{x}}$ from $\mathbf{\hat{Z}}^{p}$ and $\gamma$. The quality of the reconstructed image $\mathbf{\hat{x}}$ is evaluated using the peak signal-to-noise ratio (PSNR), which quantifies pixel-wise distortion:
${\textrm{PSNR}}:=10\log_{10}\frac{255^{2}}{{\textrm{MSE}}}~{}\text{dB},$
where MSE denotes the mean squared error between the original and reconstructed images. In this paper, our objective is to improve the PSNR in the ROI (${\textrm{PSNR}}_{{\textrm{ROI}}}$) while maintaining comparable average PSNR (${\textrm{PSNR}}_{{\textrm{Avg}}}$) with respect to previous deepJSCC models. In the following subsections, we introduce the encoder and decoder architectures of ROI-JSCC, the ROI-based loss function, and present the detailed procedure for ROI-adaptive bandwidth allocation.
II-A Encoder and Decoder Design
Figure 2: The structures of the encoder and the decoder.
The architectures of the encoder $f_{\theta}$ and decoder $g_{\phi}$ are illustrated in Figure 2. Both consist of $L$ stages. For the encoder, at stage $i$, a downsampling layer reduces the feature resolution from $\frac{H}{2^{i-1}}\times\frac{W}{2^{i-1}}\times C_{i-1}$ to $\frac{H}{2^{i}}\times\frac{W}{2^{i}}\times C_{i}$. In our model, a patch embedding layer is used at stage 1, while patch merging layers [3] are employed for the subsequent stages. Each stage then processes features using a region-of-interest-guided group (ROI group), which consists of $n_{i}$ region-of-interest-guided blocks (ROI blocks). Leveraging the ROI position $\gamma$, these ROI blocks prioritize enhancing features relevant to the ROI while still refining others. Finally, a fully connected (FC) layer maps each feature dimension from $C_{L}$ to $2C_{M}$. Through this process, the encoder extracts the semantic feature matrix $\mathbf{Z}\in\mathbb{C}^{B\times C_{m}}$ from the input image $\mathbf{x}$, where $B:=\frac{H}{2^{L}}\times\frac{W}{2^{L}}$.
For the decoder, an FC layer first transforms each feature dimension from $2C_{M}$ back to $C_{L}$. Then, for each stage $i$, an ROI group processes the features using $\gamma$, followed by a patch division layer [3] that upsamples the features from $\frac{H}{2^{L+1-i}}\times\frac{W}{2^{L+1-i}}\times C_{L+1-i}$ to $\frac{H}{2^{L-i}}\times\frac{W}{2^{L-i}}\times C_{L-i}$. Through this decoding pipeline, the decoder reconstructs the image $\mathbf{\hat{x}}$ from the input feature matrix $\mathbf{\hat{Z}}^{p}$. In the following subsection, we describe the internal structure of the ROI block in detail.
II-B ROI Block Details
Figure 3: ROI block structure and process for ROI $\gamma=(2,2)$.
Figure 4: The ROI importance mask generation process.
The detailed structure of the proposed ROI block is illustrated in Figure 3.111For clarity, the refining process is visualized using an image. However, in practice, the ROI block operates on intermediate neural features rather than raw pixel data. First, the ROI importance mask $\gamma_{m}$ is obtained through the generation process depicted in Figure 4. Based on the ROI position $\gamma$, the ROI map is constructed by labeling patches adjacent to the ROI as ROP and all other patches as RONI. Next, the ROI embedding $\gamma_{e}$ is produced by assigning the values $1.0$, $0.5$, and $0.0$ to the ROI, ROP, and RONI patches, respectively. The ROI importance mask $\gamma_{m}$ is then generated by applying the nearest-neighbor method to upsample $\gamma_{e}$ to match the resolution of the ROI block’s input feature $\mathbf{y}_{{\textrm{in}}}$. To effectively incorporate ROI information into $\mathbf{y}_{{\textrm{in}}}$ without degrading its original content, $\gamma_{m}$ is further converted into an importance feature map $\gamma_{f}$, which is then added to $\mathbf{y}_{{\textrm{in}}}$. This integration mechanism enables subsequent layers to identify and emphasize features that are most relevant for improving ${\textrm{PSNR}}_{\textrm{ROI}}$.
Then, the feature vectors in $\mathbf{y}_{{\textrm{in}}}$ are categorized based on the values in the ROI importance mask $\gamma_{m}$: features corresponding to values of $1$, $0.5$, and $0.0$ are defined as ROI feature vectors $\mathbf{z_{\textrm{ROI}}}$, ROP feature vectors $\mathbf{z_{\textrm{ROP}}}$, and RONI feature vectors $\mathbf{z_{\textrm{RONI}}}$, respectively. To enhance the processing capability of $\mathbf{z_{\textrm{ROI}}}$ without increasing computational cost, the ROI block applies separate processing strategies to $\mathbf{z_{\textrm{ROI}}}$, $\mathbf{z_{\textrm{ROP}}}$, and $\mathbf{z_{\textrm{RONI}}}$.
As illustrated in Figure 3, $\mathbf{z_{\textrm{ROI}}}$ and the majority of $\mathbf{z_{\textrm{ROP}}}$ are processed using computationally-intensive self-attention [3], whereas $\mathbf{z_{\textrm{RONI}}}$ and the $\mathbf{z_{\textrm{ROP}}}$ vectors adjacent to the three RONI are processed using lightweight spatial attention [11]. Self-attention has proven to be a powerful tool for feature extraction in deepJSCC [3], whereas spatial attention provides satisfactory performance with low computational complexity. Using both mechanisms in a manner guided by ROI, this split processing effectively enhances ${\textrm{PSNR}}_{{\textrm{ROI}}}$ while minimizing degradation in ${\textrm{PSNR}}_{{\textrm{Avg}}}$ without additional computational overhead.
In communication systems, enhancing feature correlation contributes to error correction in channel coding, while reducing redundant information is essential for efficient source coding. To achieve both objectives, the proposed ROI block processes the entire set of features jointly by applying depthwise convolution and channel attention, thereby strengthening feature correlations. In addition, it employs a gated depthwise convolutional feed-forward network (GDF) [15] to suppress redundancy and improve feature compactness.
II-C ROI-based Loss Function
Deep neural models have the capacity to learn the relationship between conditional inputs and the loss function. In our setting, the conditional input is $\gamma$, and our goal is to maximize ${\textrm{PSNR}}_{{\textrm{ROI}}}$ while minimizing the degradation in ${\textrm{PSNR}}_{{\textrm{Avg}}}$. To train ROI-JSCC for improving ${\textrm{PSNR}}_{{\textrm{ROI}}}$ across varying $\gamma$, we design the ROI-based loss function $\mathcal{L}$ as follows:
$\mathcal{L}(\mathbf{x},\mathbf{\hat{x}}):=d(\mathbf{x},\mathbf{\hat{x}})+\alpha d(\mathbf{x_{{\textrm{ROI}}}},\mathbf{\hat{x}_{{\textrm{ROI}}}})+\beta d(\mathbf{x_{{\textrm{ROP}}}},\mathbf{\hat{x}_{{\textrm{ROP}}}}),$
where $10<\beta<\alpha<$ 1 are hyperparameters and $d(\cdot)$ denotes the mean squared error (MSE). By using this loss function, ROI-JSCC learns to prioritize features that contribute to improving ${\textrm{PSNR}}_{{\textrm{ROI}}}$ with the aid of $\gamma$, while preserving ${\textrm{PSNR}}_{{\textrm{Avg}}}$. Consequently, it effectively maintains high quality in the ROI during image transmission under varying $\gamma$ without fine tuning.
II-D ROI-based Bandwidth Allocation
The possible average transmittable dimension per feature vector $C_{{\textrm{Avg}}}$ under the bandwidth constraint $k$ can be calculated by the following equation: $B\cdot C_{{\textrm{Avg}}}=k$, where $B$ is the total number of feature vectors. When the image is divided into $n_{h}\times n_{w}$ patches, the number of ROI patches $n_{\textrm{ROI}}$, ROP patches $n_{\textrm{ROP}}$, and RONI patches $n_{\textrm{RONI}}$ satisfy the following patch number inequalities: $n_{\textrm{ROP}}\leq 8n_{\textrm{ROI}}$, $n_{\textrm{RONI}}\leq n_{h}n_{w}-9n_{\textrm{ROI}}$ since each ROI patch can have at most $8$ adjacent ROP patches. Let $C_{\textrm{ROI}}$ and $C_{\textrm{RONI}}$ be the increased and decreased dimension for ROI and RONI feature vectors. Then, the bandwidth constraint $k$ is satisfied if the next inequality holds.
$n_{\textrm{ROI}}\cdot C_{{\textrm{ROI}}}\leq n_{\textrm{RONI}}\cdot C_{\textrm{RONI}}\leq(n_{h}n_{w}-9n_{\textrm{ROI}})C_{\textrm{RONI}},$
by the patch number inequalities. Therefore, with $\eta=(n_{h}n_{w}-9n_{\textrm{ROI}})/n_{\textrm{ROI}}$, assigning dimensions of $(1+\eta\tau)C_{{\textrm{Avg}}}$, $C_{{\textrm{Avg}}}$, and $(1-\tau)C_{{\textrm{Avg}}}$ to $\mathbf{z_{\textrm{ROI}}}$, $\mathbf{z_{\textrm{ROP}}}$, and $\mathbf{z_{\textrm{RONI}}}$, respectively, satisfies the bandwidth constraint $k$, where $10<\tau<$ 1 is a hyperparameter. This ROI-based bandwidth allocation significantly increases the bandwidth assigned to $\mathbf{z_{\textrm{ROI}}}$ while marginally reducing that for $\mathbf{z_{\textrm{RONI}}}$. As a result, the ROI quality is enhanced with only a minimal loss in other regions.
Refer to caption   Refer to caption
Refer to captionRefer to captionRefer to caption   Refer to caption
Figure 5: The left figures of the vertical line show ${\textrm{PSNR}}_{{\textrm{ROI}}}$ (solid lines) and ${\textrm{PSNR}}_{{\textrm{Avg}}}$ (dotted lines) results of DIV2K validation dataset under different CPP and SNR environments. The right figure of the vertical line shows the ${\textrm{PSNR}}_{{\textrm{ROI}}}$ results of the DIV2K validation dataset for the ablation study of the ROI-based mechanisms in ROI-JSCC.
Figure 6: The ROI quality of the reconstructed Kodak image at ${\textrm{CPP}}=1/12$, ${\textrm{SNR}}=4$dB. The red box is the ROI of the transmitted image.
III Experiment
III-A Experimental Setting
Baselines: To demonstrate, we compare ROI-JSCC with previous deepJSCC models ranging from the original [13] to the latest [12]. The baselines include a convolution block-based ConvJSCC [13], a residual block-based [14] ResJSCC, a Swin transformer block-based SwinJSCC [3], and a feature importance-aware block-based FAJSCC [12]. Detailed implementation codes are available at https://github.com/hansung-choi/ROI-JSCC.
Training and Testing: We use the DIV2K dataset [16] for training. The training set consists of 800 high-resolution images, randomly cropped into $256\times 256$ patches with a batch size of $11$ over $200$ epochs. The Adam optimizer with a learning rate of $0.0001$ is used to minimize the ROI-based loss function $\mathcal{L}$ (with $n_{h}=n_{w}=4$, varying $\gamma$, $\alpha=1.0$, $\beta=0.5$, and $\tau=0.1$) for ROI-JSCC, while other models are trained to minimize the MSE loss. For numerical evaluation, images of the DIV2K validation dataset are center-cropped to the nearest multiple of $128$ to match the original size closely. For visual comparison, the Kodak dataset is used.
III-B Experimental Result
In Figure 5, the three left figures of the vertical line present the ${\textrm{PSNR}}_{{\textrm{ROI}}}$ (solid lines) and ${\textrm{PSNR}}_{{\textrm{Avg}}}$ (dotted lines) results of DIV2K validation dataset across various CPP and SNR scenarios. During testing, the position of ROI $\gamma$ is randomly sampled from $41<h_{\gamma},w_{\gamma}<$ 4. As shown, our ROI-JSCC consistently achieves substantial improvements in ${\textrm{PSNR}}_{{\textrm{ROI}}}$ across all CPP and SNR conditions, outperforming other baselines. These improvements can be attributed to the four aforementioned ROI-based mechanisms, which effectively boost ${\textrm{PSNR}}_{{\textrm{ROI}}}$ while minimizing the impact on ${\textrm{PSNR}}_{{\textrm{Avg}}}$. Regarding ${\textrm{PSNR}}_{{\textrm{Avg}}}$, ROI-JSCC significantly outperforms ConvJSCC and ResJSCC, despite its focus on improving ${\textrm{PSNR}}_{{\textrm{ROI}}}$. While ROI-JSCC underperforms compared to recent state-of-the-art (SOTA) FAJSCC and SwinJSCC at low SNR, it matches or surpasses them at high SNR. This is because sacrificing more ${\textrm{PSNR}}_{{\textrm{Avg}}}$ at low SNR is necessary to improve ${\textrm{PSNR}}_{{\textrm{ROI}}}$, given that enhancing ${\textrm{PSNR}}_{{\textrm{ROI}}}$ under low SNR conditions is more challenging than at high SNR.
Figure 6 compares ROI visualization results for the reconstructed Kodak image at ${\textrm{CPP}}=\frac{1}{12}$ and ${\textrm{SNR}}=4$dB. In this result, our ROI-JSCC shows the clearest reconstruction with the highest ${\textrm{PSNR}}_{{\textrm{ROI}}}$. For instance, FAJSCC induces color jittering in the text, while SwinJSCC causes blurring near the connection between the wing and the main body. In the case of ConvJSCC, the degradation is so severe that the text becomes barely readable. In contrast, ROI-JSCC effectively reconstructs the image without these artifacts and preserves high ROI quality. In the case of ${\textrm{PSNR}}_{{\textrm{Avg}}}$, compared to the recent two SOTA models, FAJSCC ($28.01$dB) and SwinJSCC ($27.81$dB), our ROI-JSCC ($27.79$dB) shows minimal degradation. Moreover, our ROI-JSCC uses fewer computational resources ($14.13$ GFLOPs) than the recent two SOTA models, FAJSCC ($14.57$ GFLOPs) and SwinJSCC ($20.04$ GFLOPs). This efficiency is attributed to the ROI-based split processing, which enhances ${\textrm{PSNR}}_{{\textrm{ROI}}}$ without incurring additional computational overhead.
III-C Ablation Study
We conduct an ablation study to verify the effectiveness of the four ROI-based mechanisms in ROI-JSCC. In Figure 5, the right figure of the vertical line shows the ${\textrm{PSNR}}_{{\textrm{ROI}}}$ results of the DIV2K validation dataset for different configurations where some of the four ROI-based mechanisms in ROI-JSCC are added or removed. In this figure, “w/o RB,” “w/RLB,” “w/RB,” and “w/RL” denote without ROI-based bandwidth allocation, with ROI-based loss and bandwidth allocation, with ROI-based bandwidth allocation, and with ROI-based loss, respectively. Using only the ROI-based loss (FAJSCC w/RL) or only the ROI-based bandwidth allocation (FAJSCC w/RB) does not lead to improvements in ${\textrm{PSNR}}_{{\textrm{ROI}}}$. For w/RL, the ROI-based loss alone confuses FAJSCC since it cannot identify which features are related to ${\textrm{PSNR}}_{{\textrm{ROI}}}$. Similarly, for w/RB, the ROI-based bandwidth allocation also confuses FAJSCC because it does not understand why certain features are allocated different bandwidths. This indicates that simply applying these vanilla extensions is insufficient to increase ${\textrm{PSNR}}_{{\textrm{ROI}}}$.
When ROI-based loss and ROI-based bandwidth allocation are applied simultaneously (FAJSCC w/RLB), ${\textrm{PSNR}}_{{\textrm{ROI}}}$ improves because the decoder can indirectly infer which features relate to ${\textrm{PSNR}}_{{\textrm{ROI}}}$ from the varying bandwidth allocations. However, FAJSCC w/RLB still underperforms compared to all ROI-JSCC and most ROI-JSCC variants without ROI-based bandwidth allocation. This highlights the effectiveness of our proposed ROI feature map $\gamma_{f}$ extraction mechanism, which directly incorporates ROI information into subsequent layers. Moreover, comparing ROI-JSCC and ROI-JSCC w/o RB demonstrates that ROI-based bandwidth allocation significantly boosts ${\textrm{PSNR}}_{{\textrm{ROI}}}$ when combined with an appropriate ROI feature map extraction method (noting that FAJSCC w/RB alone does not significantly increase ${\textrm{PSNR}}_{{\textrm{ROI}}}$).
IV Conclusion
In this paper, we propose a region-of-interest-guided deep joint source-channel coding (ROI-JSCC) method for user experience-oriented image transmission. Unlike previous deepJSCC approaches, our ROI-JSCC adaptively preserves high quality in the region of interest (ROI), which is strongly correlated with user experience. This is achieved through ROI-based feature map extraction and a tailored loss function, allowing ROI-JSCC to learn which features are related to the ROI and how to encode them effectively for ROI quality preservation in the transmitted image. Additionally, our ROI-specific bandwidth allocation further enhances the quality of the ROI. Moreover, the ROI-based split processing improves ROI quality without introducing additional computational burden compared to two recent SOTA models. We evaluated the performance of ROI-JSCC in terms of both ROI and overall image quality, comparing it against models ranging from the original to the most recent SOTA. Furthermore, our ablation study successfully verifies the effectiveness of our proposed mechanisms in improving ROI quality.